An automotive archives gas mileage of 2.97 mi/gal. How many  kilometers can it travel on 1.5 l of gasoline?
Answer #1

Mileage = 2.9 mi/gal

1 mi = 1.609 Km

1 gal = 3.785 Liters

Now, mileage = 2.9*1.609 / 3.785 = 1.233 Km/Liter

So, travelled distance = 1.233*1.5 = 1.85 Km .... Answer